OpenStudy (anonymous):

a man walked north 119 feet and then west 120 feet.find the shortest distance between the start and end points

OpenStudy (anonymous):

you could use the pythagorean theorem

OpenStudy (anonymous):

to find the line that connects those two points

OpenStudy (anonymous):

c^2 = a^2 + b^2 c^2 = 119^2 + 120^2 c = 169